OpenStudy (anonymous):

What are the steps to solve the equation 7x − 9 = 47 for x?

OpenStudy (anonymous):

first you add nine to both sides.

OpenStudy (anonymous):

Then do I multiply or divid by 7?

OpenStudy (cj7529):

7x- 9 = 47 add 9 to both sides 7x = 47 + 9 =56 divide both sides by 7 x = 56/7 x = 8

OpenStudy (anonymous):

then you divide both side by seven x = 8

OpenStudy (anonymous):

Thank you so much! :)

OpenStudy (anonymous):

you do the opposite of what it is since it is minus 9 you add and the seven is multiplied by x so you divide to even it out and youre welcome